if cos a + cos b = a, sin a + sin b = y and theta is the arithmetic mean between a an b then sin 2theta + cos 2theta is

Grade:11

1 Answers

Nishant Vora IIT Patna
9 years ago
Hello Student,Please find the answer to your question below

Cosa + Cosb = 2Cos[(a+b)/2]Cos[(a-b)/2] = x
Sina + Sinb = 2Sin[(a+b)/2]Cos[(a-b)/2] = y
tan[(a +b )/2 ] = y/x
so tan(theta) = y/x

so now find tan(2 theta) from here once you get tan(2 theta) then you can also get sin 2theta and cos 2theta